OpenStudy (anonymous):

Solve for x: x=xe^(3-x).

OpenStudy (inkyvoyd):

take the natural log of both sides.

OpenStudy (inkyvoyd):

And, I forget log laws.

OpenStudy (anonymous):

Can anyone help me? 3.41 x 1,3019234 divided by 12347 x 8 + your age - 1804351 x 235-135425 squaredx 10 to the fifth power + 1- 1243818723847

OpenStudy (anonymous):

Post a new question instead of posting here please, Fweaa.

OpenStudy (anonymous):

\[\LARGE x=x\cdot e^{3-x}\] \[\LARGE \frac xx = e^{3-x}\] \[\LARGE 1= e^{3-x}\] \[\LARGE 1=a^x \quad \Longleftrightarrow x=0\]

OpenStudy (anonymous):

so... \[\LARGE 3-x=0\] \[\LARGE -x=0-3\] \[\LARGE -x=-3\] \[\LARGE x=3\]

OpenStudy (anonymous):

yes ... @bmp

OpenStudy (anonymous):

Ah, sorry, mate. I thought that you had ended your proof and made a minor mistake at the very end. My bad. :-)